OpenStudy (anonymous):

Which expression is equivalent to this one? –5(6 + 11) A. (–5) • 6 + 11 B. (–5) + 6 + 11 C. (–5) • 6 • 11 D. (–5) • 6 + (–5) • 11

OpenStudy (anonymous):

some one pls help will give medal

OpenStudy (anonymous):

A not 100% sure

OpenStudy (sleepyhead314):

Hello @pompom123 and Welcome to OpenStudy! :) this is using the distributive property a (b + c) = (a)*(b) + (a)*(c)

OpenStudy (anonymous):

thanks

OpenStudy (sleepyhead314):

sorry, but A is not right :/

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

ok as i sead not 100% sure

OpenStudy (mathstudent55):

The distributive property of multiplication over addition: \(a(b + c) = ab + ac\) Example: \(-2(3 + 4) = (-2) \cdot 3 + (-2) \cdot 4\)

OpenStudy (mathstudent55):

Use the rule and the example to answer your own question.

OpenStudy (anonymous):

@sleepyhead314

OpenStudy (anonymous):

were are you

OpenStudy (anonymous):

i still need an answer

OpenStudy (anonymous):

i just have four more left

OpenStudy (sleepyhead314):

hi? yes :) I think mathstudent55 was explaining it rather well :) if you'd like more I was saying the distributive property a (b + c) = (a)*(b) + (a)*(c) so you want to look for the option that is like something times something plus another something times something :) does that make more sense?

OpenStudy (sleepyhead314):

take a guess and tell me what you think it is @pompom123 :)

OpenStudy (anonymous):

nooo

OpenStudy (mathstudent55):

Follow the pattern of the example above to do your problem below.